Two planes, 2000 miles apart, fly toward each other. Their speeds differ by 35 mph. If they pass each other in 3 hours, what is thespeed of each?
Answer by ewatrrr(24785)

3(r) + 3(r-35) = 2000mi
6r = 2105
r = 2105/6 = 350.8mph the other 315.8mph